### v1

# Prompt Programming: A Platform for Dialogue-based Computational Problem Solving with Generative AI Models

Abstract.
Computing students increasingly rely on generative AI tools for programming assistance, often without formal instruction or guidance. This highlights a need to teach students how to effectively interact with AI models, particularly through natural language prompts, to generate and critically evaluate code for solving computational tasks. To address this, we developed a novel platform for prompt programming that enables authentic dialogue-based interactions, supports problems involving multiple interdependent functions, and offers on-request execution of generated code. Data analysis from over students in an introductory programming course revealed high engagement, with the majority of prompts occurring within multi-turn dialogues. Problems with multiple interdependent functions encouraged iterative refinement, with progression graphs highlighting several common strategies. Students were highly selective about the code they chose to test, suggesting that on-request execution of generated code promoted critical thinking. Given the growing importance of learning dialogue-based programming with AI, we provide this tool as a publicly accessible resource, accompanied by a corpus of programming problems for educational use.
1. Introduction
Generative AI has transformed programming and the skills required to excel in this field (Denny et al., 2024c). Students are increasingly turning to AI tools to assist with programming tasks, often without any formal guidance. This highlights a need for educators to actively teach the new skills required to use these tools effectively. Traditionally, introductory courses focused on teaching students to write code by transforming clear problem statements into syntactically and semantically correct code. However, with large language models (LLMs) now capable of solving this step, there is a growing need to emphasize reading, critically evaluating generated code, and crafting precise natural language prompts to achieve desired outcomes.
Several tools have recently been described to help students develop prompt-crafting skills. For instance, one tool expands prompts into high-level pseudocode to help students structure solutions while programming, while another converts prompts into code fragments to generate feedback for code explanation tasks (Kazemitabaar et al., 2024; Smith et al., 2024). A more direct approach is the concept of ‘Prompt Problems’ which challenge students to create single-shot prompts to solve computational tasks presented visually (Denny et al., 2024b). Building on this idea, in the current work we propose several new features that improve the value of this learning-by-prompting approach: dialogue-based interactions, support for problems involving multiple functions, and on-request execution of generated code. These novel features support more authentic interactions with LLMs, promote critical evaluation of code, and enable iterative, real-world problem-solving.
In this paper, we introduce Prompt Programming, a new publicly available platform to teach students how to craft and refine prompts for LLMs. Unlike previous single-shot approaches, it enables multi-turn conversations for iterative refinement and provides on-request execution to test code (Figure 1). The platform also supports multi-function problems, encouraging integrated problem-solving (Figure 2). Through analysis of student interaction data from a large introductory programming course, we evaluate each of these novel features, guided by the following three research questions:
RQ1: How do students’ problem solving approaches for multi-function problems differ from those for single-functions?
RQ2: How does problem complexity influence the length and message size of multi-turn conversations?
RQ3: When using on-request execution, how selective are students in deciding whether to execute generated code?
2. Related work
Generative AI for computing education. Recent advances in generative AI, in particular, code-generating large language models, have the potential to drastically change the landscape of computing and programming education (Denny et al., 2024c; Prather et al., 2023). Recent work has demonstrated this potential in numerous programming education scenarios, including repairing buggy programs (Zhang et al., 2024), enhancing programming-error-messages (Wang et al., 2024; Leinonen et al., 2023; Phung et al., 2023), providing code explanations (Sarsa et al., 2022; MacNeil et al., 2023), and generating new programming exercises (Sarsa et al., 2022; Pădurean et al., 2024) or questions (Doughty et al., 2024; Li et al., 2023). Moreover, generative AI has been leveraged to create debugging exercises that help learners identify and fix code errors (Ma et al., 2024; Pădurean et al., 2025) or for enabling conversational programming (Denny et al., 2023). It can also act as an AI pair programmer, providing real-time feedback (Lohr et al., 2025; Zamfirescu-Pereira et al., 2023) and fostering collaboration (Mozannar et al., 2024; Ma et al., 2023; Vaithilingam et al., 2022; Liu et al., 2024), or support mixed-modality tasks with both text and code (Rawal et al., 2024; Ahmed et al., 2025). These advancements highlight the potential of generative AI to address diverse educational challenges and offer unique support in novice programming environments.
Prompt problems. Denny et al. (Denny et al., 2024b) introduced the idea of ‘Prompt Problems’, a novel exercise where students craft natural language prompts to solve computational tasks presented visually. Accompanied by a prototype tool called Promptly, this approach evaluates code generated from student-written prompts against a test suite. Implemented in CS1 and CS2 courses, their study revealed positive student engagement, with learners appreciating how these exercises foster computational thinking and expose them to new constructs. However, the prototype tool presented in this work supports only single-turn prompts, which limits how well it can be applied to more complex, real-world tasks.
Kerslake et al. (Kerslake et al., 2024) investigated the integration of natural language prompting tasks into introductory programming courses, shifting the focus from syntax mastery to problem-solving. In addition to Prompt Problems, another type of task was introduced that focused on code explanation – students crafted prompts to generate code equivalent to provided code fragments. Their study found that these tasks engaged a broader range of cognitive skills compared to traditional programming assessments and appealed to students who found syntax challenging. The work highlights the potential of prompt-focused activities to reduce barriers for novices and encourage diverse participation in programming.
Most recently, Prather et al. (Prather et al., 2024a) explored the use of Prompt Problems in multilingual settings to address barriers faced by non-native English-speaking (NNES) students in programming education. The study found that NNES students successfully used their native languages for solving computational tasks through LLMs, although challenges in terminology persisted. Moreover, this work utilized the same prototype tool described by Denny et al. (Denny et al., 2024b), which was limited to single-function problems and single-prompt interactions, and is not publicly available.
Generative AI-based conversational agents for education. Generative AI conversational agents are increasingly recognized for their transformative potential across diverse educational contexts (Denny et al., 2024a). These agents are used to simulate classroom environments for teacher training (Markel et al., 2023; Lee et al., 2023), enable students to enhance their understanding by teaching AI peers (Schmucker et al., 2024), and model student behavior (Nguyen et al., 2024). They can help in boosting creativity (Acun and Acun, 2023; Bailis et al., 2023) and critical thinking (Sonkar et al., 2024) through interactive tasks, puzzles, and problem-solving, demonstrating their versatility in reshaping education.
3. Methodology
This section describes our novel Prompt Programming platform. We highlight its core functionality, then detail our classroom deployment and our approach to analyzing student interactions.
3.1. Prompt Programming Platform
We developed the Prompt Programming platform to support iterative problem-solving with LLMs, focusing on three core functionalities: multi-function problems, multi-turn conversations, and on-request code execution. These create an environment where students can iteratively refine prompts, critically evaluate model outputs prior to code execution, and develop a systematic approach to solving complex programming problems. By promoting iterative refinement of prompts and deeper engagement with code, the platform fosters critical thinking and problem-solving skills. A web app demonstrating these functionalities is available online, supporting both C and Python. We detail each of the three functionalities below.
First, the platform introduces multi-function problems, which allows students to practice solving complex tasks, mirroring real-world usage of generative models. To help students gradually adapt, we start with single-function problems, inspired by earlier works (Denny et al., 2024b), as illustrated in Figure 1(a). Building on this foundation, we introduce multi-function problems, that challenge students to guide the LLM in implementing interdependent functions and creating structures or classes. Input-output specifications are provided for each function, as illustrated in Figure 2(a), so students understand expected behaviors. A problem is considered solved only when all functions and constructs meet their respective specifications, encouraging students to ensure all components work together cohesively.
Second, the platform provides a chat-like interface for natural, multi-turn conversations, enabling students to iteratively refine their prompts, as shown in Figures 1(b) and 2(b). Each time a student sends a message, the LLM processes it along with the full conversation history to generate context-aware responses. Besides the student’s verbatim message, we prepend a system prompt specifying the programming language and instructing the LLM to exclude testing code or a main function. A ‘Reset’ button allows students to clear the current conversation history and start afresh.
Finally, the Prompt Programming platform includes on-request code execution, enabling students to critically evaluate generated code before running it against hidden test cases. Using the ‘Run Code’ button, students can execute the most recently generated block of code. A counter tracks code execution requests, which can incentivize students to be selective in their use of this feature. This functionality fosters code comprehension and critical judgment by encouraging students to hypothesize code behavior and address potential issues prior to code execution.
3.2. Study Setup
In order to investigate how our platform supports students in prompt writing and iterative problem solving, we deployed our Prompt Programming platform as part of a large introductory C programming course involving students, taught at Institution anonymized for review. The platform utilized GPT-4o mini (OpenAI, 2024), popular for its accessibility and low-cost usage. The primary goal of the deployment was to collect data on how students used features like multi-turn conversations and on-request code execution to tackle progressively complex problems, including multi-function problems that mirror real-world programming scenarios. To evaluate their engagement and usage patterns, we collected interaction data, including student queries, the LLM’s responses, and code execution requests. Students were also invited to provide optional open-response feedback, providing qualitative insights into their interactions with Prompt Programming.
The study spanned three laboratory assignments (L7, L9, and L10), each lasting one week and each comprising problems created by the authors, who have extensive teaching experience in programming. The problems were chosen to align with the topics students were learning during the course at the time of each laboratory. In L7, students needed to solve all simple single-function problems: counting negatives in an array (see Figure 1), summing even numbers in an array, and finding the index of the last zero in an array. In L9, students needed to solve at least of the more complex single-function problems: sorting a subarray within specified indices, updating a matrix by propagating s across rows and columns, and adding binary numbers represented as arrays. Finally, in L10, students needed to solve at least of the multi-function problems, which required helper functions and structures: password validation (see Figure 2), guiding a robot’s movement on a grid using a structure and functions for turning and advancing, and checking if an expression’s round parentheses are balanced using a stack and corresponding functions. To prevent students from getting sidetracked within a single conversation, we limited conversations to messages for L7 and for L9 and L10 (conversations would need to be reset by the student once reaching these limits).
4. Results
Next, we summarize participation, present results for each research question (RQ), and highlight key insights from student reflections.
4.1. Descriptive Statistics
We begin by reporting the success rate of students across three labs. Recall that L7 required solving out of problems, while L9 and L10 required solving only out of problems. Out of students in the course, typically about of students participate in a given laboratory assignment. For students who participated in a given lab, the success rates are as follows: for L7, for L9, and for L10. The average number of conversations and average total messages by a student for a problem in a given lab are: (i) and for L7, (ii) and for L9, and (iii) and for L10. Interestingly, the success rates are high even for difficult problems, though the average number of messages increased with lab difficulty.
4.2. RQ1: Solving Multi-function Problems
To understand how students approach multi-function problems, we analyzed the progression of correctly implemented functions in conversations from L10. We took the first successful conversation for each student (i.e., the first conversation where they managed to solve the problem) and modeled progression as a directed graph, where nodes represent cumulative states of correctly implemented functions. A transition was recorded as an edge whenever new functions were added to the set of correctly implemented functions. We calculated edge weights based on the number of students whose conversations included a specific transition, highlighting the most frequent paths. To simplify visualization, we filtered the graph to include the top edges in terms of weights, focusing on the most common transitions students followed to solve multi-function problems.
Figures 3(a) and 3(b) show that many students attempted to implement all functions in one step. However, the figures also highlight a large diversity of paths, with students using multi-turn conversations to iteratively refine functions and address errors while retaining correct components. Both problems revealed a hub-like intermediate state, where transitions were concentrated. This state often marked a point where most helper functions were correctly implemented, providing a foundation for tackling advanced aspects of the problem. These findings highlight that multi-function problems do encourage diverse problem-solving strategies – in Section 4.6 we discuss how support for the stepwise development of solutions could be better enabled in future work.
4.3. RQ2: Usage of Multi-turn Conversations
To investigate RQ2, we analyzed how multi-turn conversations supported students in solving problems of varying complexity. Specifically, we examined the total number of messages sent in their first successful conversation and the median character length of each message by its position (k-th message) within that conversation. This analysis provided insights into how students utilized multi-turn interactions to refine their prompts and achieve success.
Figure 4(a) highlights that students actively leveraged multi-turn conversations across all labs. In L7, most students succeeded within to messages, reflecting the simplicity of the problems. In L9, and more markedly in L10, as problem complexity increased, students made greater use of the multi-turn feature, with more than 25% of the students who solved a problem in L10 requiring more than 5 messages to refine their solutions. This demonstrates that students did not rely on single, monolithic prompts to solve the problems but instead utilized smaller, iterative steps – a more authentic and practical way to interact with generative AI tools.
Figure 4(b) further shows that the first message was generally the longest, especially in L10, where problems required detailed framing and initial setup. Subsequent messages were shorter and more consistent in length, reflecting students’ focused adjustments and iterative refinements as they progressed toward a solution. These patterns highlight the value of multi-turn conversations, enabling students to approach problem-solving incrementally and adapt their strategies based on model feedback, closely mirroring real-world interactions with AI systems.
4.4. RQ3: Usage of On-request Execution
To evaluate the on-request code execution feature, we wish to understand how selective students are in deciding whether to execute the generated code. Simply put, we analyzed whether students were more likely to execute code generated by the model when the code was correct compared to when the code was incorrect. For every student prompt that resulted in code generation, we determined the correctness of that code and then analyzed whether or not the student chose to execute it.
Figure 5 shows that across all messages that resulted in generated code, roughly - of the time code execution was requested. This provides some insight into how frequently students relied on running code to support their problem-solving process. Moreover, students were much more likely to execute correct code than incorrect code, indicating that they applied judgment in assessing whether the generated code was ready for testing. This behavior demonstrates that on-request execution does encourage students to critically evaluate the output of the model before acting, fostering thoughtful and iterative problem-solving. Importantly, students consistently exhibited this judgment across all lab sessions.
4.5. Student Reflections
To investigate student engagement with Prompt Programming, we also gathered quantitative reflections. A common instructor concern relating to generative AI (GenAI) is the risk of student over-reliance (Prather et al., 2023; Sheard et al., 2024; Prather et al., 2024b). There may be a perception that platforms like ours, which support prompt programming, could reinforce this concern by leading students to believe that all problems can be solved through natural language prompting alone. This risk could be particularly pronounced if students were only exposed to trivial problems. However, we hypothesize that as problem complexity increases, it becomes more difficult to solve tasks solely through natural language prompting, even with multi-turn conversations.
To explore this, we asked students to indicate what they felt was the ideal combination of skills in natural language (NL) prompting and manual code editing for solving harder problems. The results, shown in Figure 6, reveal a well-balanced response: the vast majority of students believed that an equal mix of these skills is most desirable. At the same time, students reported better understanding of their own code compared to AI-generated solutions, suggesting that working with AI-generated code requires additional effort to interpret and refine. Finally, students found NL prompting easier than direct coding, reflecting its ability to reduce the focus on syntax and low-level implementation details. These observations suggest teaching strategies should evolve to prioritize a hybrid approach, combining NL prompting and traditional programming skills.
4.6. Limitations
Some limitations in our study suggest avenues for future work. First, the testing framework for each multi-function problem currently uses a single driver file that validates all required functions together. This leads to a compilation error if students attempt to run the code before all required functions have been generated through (iterative) prompting. Instead, one could check each component independently, allowing for more modular testing. Second, we did not compare multi-turn and single-turn conversations through an A/B study, which would reveal insights into how each influences success and students’ problem-solving strategies. Finally, we did not analyze how students’ approaches and success correlate with their overall academic performance, leaving this relationship to be explored in future research.
5. Concluding Remarks
We introduced Prompt Programming, a dialogue-based platform designed to help students develop the skill of crafting effective prompts while supporting iterative problem-solving with generative AI. Findings from its first use in a large introductory course demonstrate that students effectively used multi-turn conversations to refine solutions, particularly for complex multi-function problems. Students also exhibited judgment when using on-request code execution, running correct code more often than erroneous code. This highlights the platform’s role in fostering skills in prompt writing, critical evaluation, and systematic debugging. Despite the focus on prompting, student reflections showed that they recognized the importance of integrating natural language prompting alongside more traditional coding skills when solving complex problems.
Prompt Programming is publicly available along with a curated set of programming exercises designed for educational use. The platform offers learners an accessible and interactive environment to practice prompt writing and iterative problem-solving in natural language. Future work will focus on expanding the platform’s capabilities by introducing direct code editing features and broadening the variety of exercises to better support diverse learning needs and skill levels. We also aim to explore multi-language interactions and measure the impact of Prompt Programming on student learning outcomes. In the long term, we plan to integrate mechanisms enabling learners to create and share new exercises, fostering a collaborative learning community.
Acknowledgements.
Funded/Cofunded by the European Union (ERC, TOPS, 101039090). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.References
- (1)
- Acun and Acun (2023) Cagla Acun and Ramazan Acun. 2023. GAI-Enhanced Assignment Framework: A Case Study on Generative AI Powered History Education. In NeurIPS’23 Workshop on Generative AI for Education.
- Ahmed et al. (2025) Umair Z Ahmed, Shubham Sahai, Ben Leong, and Amey Karkare. 2025. Feasibility Study of Augmenting Teaching Assistants with AI for CS1 Programming Feedback. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Bailis et al. (2023) Suma Bailis, Lara McConnaughey, Jane Friedhoff, Feiyang Chen, Chase Adams, and Jacob Moon. 2023. WordPlay: An Agent Framework for Language Learning Games. In NeurIPS’23 Workshop on Generative AI for Education.
- Denny et al. (2024a) Paul Denny, Sumit Gulwani, Neil T. Heffernan, Tanja Käser, Steven Moore, Anna N. Rafferty, and Adish Singla. 2024a. Generative AI for Education (GAIED): Advances, Opportunities, and Challenges. CoRR abs/2402.01580 (2024).
- Denny et al. (2023) Paul Denny, Viraj Kumar, and Nasser Giacaman. 2023. Conversing with Copilot: Exploring Prompt Engineering for Solving CS1 Problems Using Natural Language. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Denny et al. (2024b) Paul Denny, Juho Leinonen, James Prather, Andrew Luxton-Reilly, Thezyrie Amarouche, Brett A. Becker, and Brent N. Reeves. 2024b. Prompt Problems: A New Programming Exercise for the Generative AI Era. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Denny et al. (2024c) Paul Denny, James Prather, Brett A. Becker, James Finnie-Ansley, Arto Hellas, Juho Leinonen, Andrew Luxton-Reilly, Brent N. Reeves, Eddie Antonio Santos, and Sami Sarsa. 2024c. Computing Education in the Era of Generative AI. Commun. ACM (2024).
- Doughty et al. (2024) Jacob Doughty et al. 2024. A Comparative Study of AI-Generated (GPT-4) and Human-crafted MCQs in Programming Education. In Proceedings of the Australasian Computing Education Conference (ACE).
- Kazemitabaar et al. (2024) Majeed Kazemitabaar, Runlong Ye, Xiaoning Wang, Austin Zachary Henley, Paul Denny, Michelle Craig, and Tovi Grossman. 2024. CodeAid: Evaluating a Classroom Deployment of an LLM-based Programming Assistant that Balances Student and Educator Needs. In Proceedings of the Conference on Human Factors in Computing Systems (CHI).
- Kerslake et al. (2024) Chris Kerslake, Paul Denny, David H. Smith, James Prather, Juho Leinonen, Andrew Luxton-Reilly, and Stephen MacNeil. 2024. Integrating Natural Language Prompting Tasks in Introductory Programming Courses. In Proceedings of the Virtual Global Computing Education Conference (SIGCSE Virtual).
- Lee et al. (2023) Unggi Lee et al. 2023. Generative Agent for Teacher Training: Designing Educational Problem-Solving Simulations with Large Language Model-based Agents for Pre-Service Teachers. In NeurIPS’23 Workshop on Generative AI for Education.
- Leinonen et al. (2023) Juho Leinonen, Arto Hellas, Sami Sarsa, Brent N. Reeves, Paul Denny, James Prather, and Brett A. Becker. 2023. Using Large Language Models to Enhance Programming Error Messages. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Li et al. (2023) Yi Li, Riteng Zhang, Danni Qu, and Maíra Marques Samary. 2023. Leveraging LLMs and MLPs in Designing a Computer Science Placement Test System. In Proceedings of the International Conference on Computational Science and Computational Intelligence (CSCI).
- Liu et al. (2024) Suqing Liu, Zezhu Yu, Feiran Huang, Yousef Bulbulia, Andreas Bergen, and Michael Liut. 2024. Can Small Language Models With Retrieval-Augmented Generation Replace Large Language Models When Learning Computer Science?. In Proceedings of the Conference on Innovation and Technology in Computer Science Education (ItiCSE).
- Lohr et al. (2025) Dominic Lohr, Hieke Keuning, and Natalie Kiesler. 2025. You’re (Not) My Type - Can LLMs Generate Feedback of Specific Types for Introductory Programming Tasks? Journal of Computer Assisted Learning (2025).
- Ma et al. (2024) Qianou Ma, Hua Shen, Kenneth Koedinger, and Sherry Tongshuang Wu. 2024. How to Teach Programming in the AI Era? Using LLMs as a Teachable Agent for Debugging. In Proceeding of the International Conference on Artificial Intelligence in Education (AIED).
- Ma et al. (2023) Qianou Christina Ma, Sherry Tongshuang Wu, and Ken Koedinger. 2023. Is AI the Better Programming Partner? Human-Human Pair Programming vs. Human-AI pAIr Programming. In AIED Workshop on Empowering Education with LLMs.
- MacNeil et al. (2023) Stephen MacNeil, Andrew Tran, Arto Hellas, Joanne Kim, Sami Sarsa, Paul Denny, Seth Bernstein, and Juho Leinonen. 2023. Experiences from Using Code Explanations Generated by Large Language Models in a Web Software Development E-Book. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Markel et al. (2023) Julia M. Markel, Steven G. Opferman, James A. Landay, and Chris Piech. 2023. GPTeach: Interactive TA Training with GPT-based Students. In Proceedings of the Conference on Learning @ Scale (L@S).
- Mozannar et al. (2024) Hussein Mozannar, Gagan Bansal, Adam Fourney, and Eric Horvitz. 2024. Reading Between the Lines: Modeling User Behavior and Costs in AI-Assisted Programming. In Proceedings of the Conference on Human Factors in Computing Systems (CHI).
- Nguyen et al. (2024) Manh Hung Nguyen, Sebastian Tschiatschek, and Adish Singla. 2024. Large Language Models for In-Context Student Modeling: Synthesizing Student’s Behavior in Visual Programming from One-Shot Observation. In Proceedings of the International Conference on Educational Data Mining (EDM).
- OpenAI (2024) OpenAI. 2024. GPT-4o mini: Advancing Cost-efficient Intelligence. https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/.
- Pădurean et al. (2025) Victor-Alexandru Pădurean, Paul Denny, and Adish Singla. 2025. BugSpotter: Automated Generation of Code Debugging Exercises. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Pădurean et al. (2024) Victor-Alexandru Pădurean, Georgios Tzannetos, and Adish Singla. 2024. Neural Task Synthesis for Visual Programming. Transactions on Machine Learning Research (TMLR) (2024).
- Phung et al. (2023) Tung Phung, José Cambronero, Sumit Gulwani, Tobias Kohn, Rupak Majumdar, Adish Singla, and Gustavo Soares. 2023. Generating High-Precision Feedback for Programming Syntax Errors using Large Language Models. In Proceedings of the International Conference on Educational Data Mining (EDM).
- Prather et al. (2023) James Prather et al. 2023. The Robots Are Here: Navigating the Generative AI Revolution in Computing Education. In Proceedings of the Working Group Reports of the Conference on Innovation and Technology in Computer Science Education (ItiCSE).
- Prather et al. (2024a) James Prather, Brent N. Reeves, Paul Denny, Juho Leinonen, Stephen MacNeil, Andrew Luxton-Reilly, João Orvalho, Amin Alipour, Ali Alfageeh, Thezyrie Amarouche, Bailey Kimmel, Jared Wright, Musa Blake, and Gweneth Barbre. 2024a. Breaking the Programming Language Barrier: Multilingual Prompting to Empower Non-Native English Learners. CoRR abs/2412.12800 (2024).
- Prather et al. (2024b) James Prather, Brent N. Reeves, Juho Leinonen, Stephen MacNeil, Arisoa S. Randrianasolo, Brett A. Becker, Bailey Kimmel, Jared Wright, and Ben Briggs. 2024b. The Widening Gap: The Benefits and Harms of Generative AI for Novice Programmers. In Proceedings of the Conference on International Computing Education Research (ICER).
- Rawal et al. (2024) Ruchit Rawal, Victor-Alexandru Pădurean, Sven Apel, Adish Singla, and Mariya Toneva. 2024. Hints Help Finding and Fixing Bugs Differently in Python and Text-based Program Representations. CoRR abs/2412.12471 (2024).
- Sarsa et al. (2022) Sami Sarsa, Paul Denny, Arto Hellas, and Juho Leinonen. 2022. Automatic Generation of Programming Exercises and Code Explanations Using Large Language Models. In Proceedings of the Conference on International Computing Education Research (ICER).
- Schmucker et al. (2024) Robin Schmucker, Meng Xia, Amos Azaria, and Tom M. Mitchell. 2024. Ruffle &Riley: Insights from Designing and Evaluating a Large Language Model-Based Conversational Tutoring System. In Proceeding of the International Conference on Artificial Intelligence in Education (AIED).
- Sheard et al. (2024) Judy Sheard, Paul Denny, Arto Hellas, Juho Leinonen, Lauri Malmi, and Simon. 2024. Instructor Perceptions of AI Code Generation Tools - A Multi-Institutional Interview Study. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Smith et al. (2024) David H. Smith, Paul Denny, and Max Fowler. 2024. Prompting for Comprehension: Exploring the Intersection of Explain in Plain English Questions and Prompt Writing. In Proceedings of the Conference on Learning @ Scale (L@S).
- Sonkar et al. (2024) Shashank Sonkar, Xinghe Chen, Myco Le, Naiming Liu, Debshila Basu Mallick, and Richard G. Baraniuk. 2024. Code Soliloquies for Accurate Calculations in Large Language Models. In Proceedings of the Learning Analytics and Knowledge Conference (LAK).
- Vaithilingam et al. (2022) Priyan Vaithilingam, Tianyi Zhang, and Elena L. Glassman. 2022. Expectation vs. Experience: Evaluating the Usability of Code Generation Tools Powered by Large Language Models. In Proceedings of the Conference on Human Factors in Computing Systems (CHI).
- Wang et al. (2024) Sierra Wang, John C. Mitchell, and Chris Piech. 2024. A Large Scale RCT on Effective Error Messages in CS1. In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
- Zamfirescu-Pereira et al. (2023) J.D. Zamfirescu-Pereira, Laryn Qi, Bjorn Hartmann, John DeNero, and Narges Norouzi. 2023. Conversational Programming with LLM-Powered Interactive Support in an Introductory Computer Science Course. In NeurIPS’23 Workshop on Generative AI for Education.
- Zhang et al. (2024) Jialu Zhang, José Pablo Cambronero, Sumit Gulwani, Vu Le, Ruzica Piskac, Gustavo Soares, and Gust Verbruggen. 2024. PyDex: Repairing Bugs in Introductory Python Assignments using LLMs. Proceedings of the ACM on Programming Languages (2024).

### v2

@@
 Generative AI has transformed programming and the skills required to excel in this field (Denny et al., 2024c).
-Students are increasingly turning to AI tools to assist with programming tasks, often without any formal guidance.
+Students are increasingly turning to AI tools to assist with programming tasks, often without any formal guidance (Amoozadeh et al., 2024).
 This highlights a need for educators to actively teach the new skills required to use these tools effectively.
@@
 However, with large language models (LLMs) now capable of solving this step, there is a growing need to emphasize reading, critically evaluating generated code, and crafting precise natural language prompts to achieve desired outcomes.
-Several tools have recently been described to help students develop prompt-crafting skills.
-For instance, one tool expands prompts into high-level pseudocode to help students structure solutions while programming, while another converts prompts into code fragments to generate feedback for code explanation tasks (Kazemitabaar et al., 2024; Smith et al., 2024).
+Several tools were recently described to help students develop prompt-crafting skills.
+For instance, CodeAid expands prompts into high-level pseudocode to help students structure solutions while programming (Kazemitabaar et al., 2024), whereas an EiPE-inspired tool converts prompts into code fragments to generate feedback for code explanation tasks (Smith et al., 2024).
 A more direct approach is the concept of ‘Prompt Problems’ which challenge students to create single-shot prompts to solve computational tasks presented visually (Denny et al., 2024b).
@@
 These novel features support more authentic interactions with LLMs, promote critical evaluation of code, and enable iterative, real-world problem-solving.
-In this paper, we introduce Prompt Programming, a new publicly available platform to teach students how to craft and refine prompts for LLMs.
-Unlike previous single-shot approaches, it enables multi-turn conversations for iterative refinement and provides on-request execution to test code (Figure 1).
-The platform also supports multi-function problems, encouraging integrated problem-solving (Figure 2).
-Through analysis of student interaction data from a large introductory programming course, we evaluate each of these novel features, guided by the following three research questions:
-RQ1: How do students’ problem solving approaches for multi-function problems differ from those for single-functions?
+In this paper, we introduce Prompt Programming, a free web-based platform for students to practice crafting and refining prompts for LLMs, available at https://www.promptprogram.org with support for Python and C-based exercises.
+Unlike previous single-shot approaches, it enables multi-turn conversations for iterative refinement, provides on-request execution to test code (Figure 1), and supports multi-function problems (Figure 2).
+We evaluate these novel features using student data from a large introductory programming course, guided by three research questions:
+RQ1: What approaches do students employ when solving multi-function problems?
 RQ2: How does problem complexity influence the length and message size of multi-turn conversations?
@@
 Implemented in CS1 and CS2 courses, their study revealed positive student engagement, with learners appreciating how these exercises foster computational thinking and expose them to new constructs.
-However, the prototype tool presented in this work supports only single-turn prompts, which limits how well it can be applied to more complex, real-world tasks.
+However, the prototype tool presented in this prior work supports only single-turn prompts, which limits how well it can be applied to more complex, real-world tasks.
 Kerslake et al.
@@
 We developed the Prompt Programming platform to support iterative problem-solving with LLMs, focusing on three core functionalities: multi-function problems, multi-turn conversations, and on-request code execution.
-These create an environment where students can iteratively refine prompts, critically evaluate model outputs prior to code execution, and develop a systematic approach to solving complex programming problems.
+These allow students to iteratively refine prompts, critically evaluate model outputs before code execution, and develop a systematic approach to solving complex programming problems.
 By promoting iterative refinement of prompts and deeper engagement with code, the platform fosters critical thinking and problem-solving skills.
-A web app demonstrating these functionalities is available online, supporting both C and Python.
-We detail each of the three functionalities below.
+We detail each functionality below.
 First, the platform introduces multi-function problems, which allows students to practice solving complex tasks, mirroring real-world usage of generative models.
@@
 Second, the platform provides a chat-like interface for natural, multi-turn conversations, enabling students to iteratively refine their prompts, as shown in Figures 1(b) and 2(b).
-Each time a student sends a message, the LLM processes it along with the full conversation history to generate context-aware responses.
-Besides the student’s verbatim message, we prepend a system prompt specifying the programming language and instructing the LLM to exclude testing code or a main function.
-A ‘Reset’ button allows students to clear the current conversation history and start afresh.
+Each message is processed with the full conversation history to generate context-aware responses.
+Besides the student’s verbatim message, we prepend a system prompt specifying the programming language and instructing the LLM to exclude test code or a main function.
+A ‘Reset’ button lets students clear the conversation history and start afresh.
 Finally, the Prompt Programming platform includes on-request code execution, enabling students to critically evaluate generated code before running it against hidden test cases.
-Using the ‘Run Code’ button, students can execute the most recently generated block of code.
-A counter tracks code execution requests, which can incentivize students to be selective in their use of this feature.
-This functionality fosters code comprehension and critical judgment by encouraging students to hypothesize code behavior and address potential issues prior to code execution.
+With the ‘Run Code’ button, students can execute the most recently generated block of code.
+A counter tracks code execution requests, incentivizing students to be selective in using this feature.
+This functionality fosters code comprehension and critical judgment by encouraging students to hypothesize code behavior and address issues before running it.
 3.2. Study Setup
-In order to investigate how our platform supports students in prompt writing and iterative problem solving, we deployed our Prompt Programming platform as part of a large introductory C programming course involving students, taught at Institution anonymized for review.
-The platform utilized GPT-4o mini (OpenAI, 2024), popular for its accessibility and low-cost usage.
+In order to investigate how our platform supports students in prompt writing and iterative problem solving, we deployed a version of our Prompt Programming platform, specifically focusing on C exercises, in a large introductory C programming course taught at the University of Auckland with students.
+The platform used GPT-4o mini (OpenAI, 2024), popular for its accessibility and low cost.
 The primary goal of the deployment was to collect data on how students used features like multi-turn conversations and on-request code execution to tackle progressively complex problems, including multi-function problems that mirror real-world programming scenarios.
@@
 Finally, in L10, students needed to solve at least of the multi-function problems, which required helper functions and structures: password validation (see Figure 2), guiding a robot’s movement on a grid using a structure and functions for turning and advancing, and checking if an expression’s round parentheses are balanced using a stack and corresponding functions.
-To prevent students from getting sidetracked within a single conversation, we limited conversations to messages for L7 and for L9 and L10 (conversations would need to be reset by the student once reaching these limits).
+To prevent students from getting sidetracked within a single conversation, we limited conversations to messages for L7, and for L9 and L10 (conversations would need to be reset by the student once reaching these limits).
 4. Results
@@
 In L9, and more markedly in L10, as problem complexity increased, students made greater use of the multi-turn feature, with more than 25% of the students who solved a problem in L10 requiring more than 5 messages to refine their solutions.
-This demonstrates that students did not rely on single, monolithic prompts to solve the problems but instead utilized smaller, iterative steps – a more authentic and practical way to interact with generative AI tools.
+This demonstrates that students did not rely on single, monolithic prompts to solve the problems but instead utilized smaller, iterative steps – a more authentic and practical way to interact with generative AI tools (Vaithilingam et al., 2022; Denny et al., 2023).
 Figure 4(b) further shows that the first message was generally the longest, especially in L10, where problems required detailed framing and initial setup.
@@
 Some limitations in our study suggest avenues for future work.
-First, the testing framework for each multi-function problem currently uses a single driver file that validates all required functions together.
-This leads to a compilation error if students attempt to run the code before all required functions have been generated through (iterative) prompting.
-Instead, one could check each component independently, allowing for more modular testing.
-Second, we did not compare multi-turn and single-turn conversations through an A/B study, which would reveal insights into how each influences success and students’ problem-solving strategies.
-Finally, we did not analyze how students’ approaches and success correlate with their overall academic performance, leaving this relationship to be explored in future research.
+First, the testing framework for multi-function problems uses a single driver file that validates all functions together, which can lead to compilation errors if students run code before generating all required components.
+A more modular approach could enable independent validation of partial solutions.
+Second, prior exposure to L7 and L9 may have influenced student behavior in L10, such as restart frequency or message refinement.
+Third, we did not conduct an A/B study comparing single-turn and multi-turn conversations to isolate their impact on problem-solving.
+Finally, we did not examine how students’ approaches and success relate to academic performance or how reliance on LLM-generated code affects students’ long-term ability to independently generate and understand code.
+Each of these limitations highlights a direction for future investigation.
 5. Concluding Remarks
@@
 Prompt Programming is publicly available along with a curated set of programming exercises designed for educational use.
-The platform offers learners an accessible and interactive environment to practice prompt writing and iterative problem-solving in natural language.
-Future work will focus on expanding the platform’s capabilities by introducing direct code editing features and broadening the variety of exercises to better support diverse learning needs and skill levels.
+The platform offers learners an accessible, interactive environment to practice prompt writing and iterative problem-solving in natural language.
+Future work will expand its capabilities by introducing direct code editing and broadening the variety of exercises.
 We also aim to explore multi-language interactions and measure the impact of Prompt Programming on student learning outcomes.
-In the long term, we plan to integrate mechanisms enabling learners to create and share new exercises, fostering a collaborative learning community.
+In the longer term, we aim to support educators to customize and assign exercises, and to enable learners to create and share their own problems, fostering a collaborative, community-driven learning environment.
 Acknowledgements.
@@
 In Proceedings of the Technical Symposium on Computer Science Education (SIGCSE).
+- Amoozadeh et al.
+(2024) Matin Amoozadeh, Daye Nam, Daniel Prol, Ali Alfageeh, James Prather, Michael Hilton, Sruti Srinivasa Ragavan, and Amin Alipour.
+2024.
+Student-AI Interaction: A Case Study of CS1 Students.
+In Koli Calling International Conference on Computing Education Research (Koli Calling).
 - Bailis et al.
@@
 - Doughty et al.
-(2024) Jacob Doughty et al.
+(2024) Jacob Doughty, Zipiao Wan, Anishka Bompelli, Jubahed Qayum, Taozhi Wang, Juran Zhang, Yujia Zheng, Aidan Doyle, Pragnya Sridhar, Arav Agarwal, Christopher Bogart, Eric Keylor, Can Kültür, Jaromír Savelka, and Majd Sakr.
 2024.
@@
 - Lee et al.
-(2023) Unggi Lee et al.
+(2023) Unggi Lee, Sanghyeok Lee, Junbo Koh, Yeil Jeong, Haewon Jung, Gyuri Byun, Yunseo Lee, Jewoong Moon, Jieun Lim, and Hyeoncheol Kim.
 2023.
@@
 - Prather et al.
-(2023) James Prather et al.
+(2023) James Prather, Paul Denny, Juho Leinonen, Brett A.
+Becker, Ibrahim Albluwi, Michelle Craig, Hieke Keuning, Natalie Kiesler, Tobias Kohn, Andrew Luxton-Reilly, Stephen MacNeil, Andrew Petersen, Raymond Pettit, Brent N.
+Reeves, and Jaromír Savelka.
 2023.